Operations Report – May 2nd

Crew 261 Operations Report 02-05-2023
SOL: 2
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
– Power system
– Robotic observatory
Notes on non-nominal systems:
Power system: One of the 3 PV strings breaker was found disconnected at 14:30. It is unknown how long this breaker has been disconnected, but a photo showing it connected was taken on April 30 at 13:27 local time. The related MPPT tracker was indicating 0 amps and the fault LED was illuminated. As a consequence, 1/3 of the power was unavailable, and the base was running on 2/3 of the normal PV power. This was reported to Mission Support who promptly reconnected the system. The Crew Engineer will inspect the power system monitoring logs to find out how long this breaker has been disconnected and if a root cause can be determined.
Spirit rover used: Yes
Hours: (before EVA) 220.1
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 87%
Currently charging: Yes
Opportunity rover used: No
Hours: 122.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Curiosity rover used: No
Hours: 232.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 262.7
Beginning charge: 100%
Ending charge: 83%
Currently charging: Yes
General notes on rovers: Based on 2 EVAs covering a 1.2km drive each, as noted in yesterday’s report, there are large discrepancies in battery capacity/consumption (in response to yesterday’s report, Dr. Rupert mentioned Percy has older batteries, explaining a lower usable capacity).
Today, we prepared a preliminary usable battery capacity estimate based on an average consumption of 25kWh/100km. This consumption was selected as a preliminary baseline because this is Opportunity’s measured consumption (the best of the 4 rovers), and it is a usual value for sedan/SUV EVs, which comforts the relevance of this value. Preliminary usable battery capacity estimate: Opportunity: ~15kWh (design capacity is 14.9kWh), Curiosity: 5kWh, Spirit: 2.5kWh, Perseverance: 2kWh.
This being preliminary, we will refine these values over time and report.
Summary of Hab operations: Nominal
WATER USE: rate of 33 gallons per day. Cumulative use since Sol 1: 140 Gal (but a suspected faucet leak and the cylindrical shape of the tank contributed to 100 indicated gallons to be lost on Sol 1). At this rate and with today’s water usable volume available, we estimate we will be 2 Sols short, or 68 gallons short.
Water (static tank): 410 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: Nominal
Summary of suits and radios: Nominal. We tested and are now using a satellite positioning and communication system to track EVAs. The primary goal is to improve safety of the crew while on EVA, in two ways. First, a crew member in EVA can now request a rescue (to official rescue services) if a real emergency situation arises by the press of a button. Second, a crew member in EVA can now request support from other crew (in another EVA or at the base) or from mission support at the press of a few buttons. The secondary (but very useful on a daily basis) goal of this system is to enable supporting crew at the Hab to monitor EVA crew in near-real-time (2-min delay at most) on a satellite map, providing relevant context to radio comms. Furthermore, EVA crew can now mark points of interest and send those in near-real-time (<1 min delay) to said map, to add relevant context when discussing a particular location. We will continue to test this system, which will eventually be replaced by the Meshstatic infrastructure.
Summary of GreenHab operations:
WATER USE: 17.8 gallons (Daily water usage for crops: 8,5 gal at 06:40 AM + 6,3 gallons at 07:00 PM. Total = 14,8 gal / Daily water usage for research and/or other purposes: 3 gal.)
Heater: Off
Supplemental light: Off
Harvest: Cucumbers: 1223g, mint: 23g
Summary of ScienceDome operations: Spirulina bioreactor device installation finalized. Culture media preparation completed (8 liters were produced). Culture has been received this evening by cargo. Culture has been partially inserted into the bioreactor (5 liters of media) due to 3 remaining liters of media being still too hot. As a result, 2/3 of the first batch of spirulina culture has been inserted. 1/3 of the first batch is ready for insertion tomorrow. The second batch is standing by as backup. Computer setup is completed and needs to be refined over the next few days to ensure seamless operation (currently, as power loss/restart will not automatically start the experiment).
Biostimulant solution was prepared using spirulina and compost.
Dual split: Off
Summary of RAM operations: (Tools used, work done) Our Crew Robotics Engineer made excellent progress in the assembly of the Atmosphinder robot, using own tools. Assembly is 80% complete. Expecting 100% tomorrow and structural tests.
Summary of any observatory issues: Solar observations have been made, images processed and one stellar image has been produced.
Summary of health and safety issues: Nominal day
Questions, concerns and requests to Mission Support: Thank you for your support throughout the day.

Copyright © The Mars Society. All rights reserved. | Main Site